OpenStudy (owlcoffee):

So I have encountered during a problem-solving excercise a function with the form: \[h(x)= 3e ^{3x}-8x\] And I am interested to see if this is solvable algebraically, which I tried and got the solution \(x=\frac{ W(-\frac{ 9 }{ 8 }) }{ -3 }\) wanted to check if I got it right and would gladly appreciate if I got some feedback.

OpenStudy (whpalmer4):

That is correct, and you can read more about the function here: http://mathworld.wolfram.com/LambertW-Function.html

OpenStudy (kainui):

One way would be to plug it back in: \[3e^{-W(-9/8)} + \frac{8}{3} W(-9/8)\] Then simplify that left term with this rearrangement: \[W(x)e^{W(x)}=x\] into \[\frac{W(x)}{x}=e^{-W(x)}\] that'll give you \[3\frac{8}{9}W(-9/8) + \frac{8}{3} W(-9/8) \] Which clearly equals 0 (I assume you were looking for the roots of h(x) I guess?)

OpenStudy (owlcoffee):

Yes, So I got the correct answer. How would you go about approximating the value of the solution using Newtons Method?

OpenStudy (kainui):

Is that what your assignment is to do, or is this what you believe is the best way to evaluate the product log?

OpenStudy (owlcoffee):

Oh, I thought it could've been the best way to approximate it, but is there one which is not possibly restricted (Newtons method sometimes does fail).

OpenStudy (kainui):

Depends, like I'm asking, what's the purpose here. If it was up to me, I would just throw it in maple or mathematica and be done with it at this point if I needed some actual physical approximation to something. Otherwise leaving it like this is just as well as \(\sqrt{ \ln x } \) to me.

OpenStudy (kainui):

One cheap trick to evaluating it could be \[W(x)e^{W(x)} = x\] take the logarithm: \[\ln W(x) + W(x) = \ln x \] Now you have a recursive way to calculate it simply that's dead easy to implement with no trouble: \[W(x) = \ln x - \ln W(x)\] At least that's what I've done in the past, I think just using Newton's method just like you would normally should work too, I think I have some papers on evaluating it actually with other numerical methods if I look. Here's another fun problem to try to evaluate exactly while we're on the subject: \[W(-\tfrac{\pi}{2})\]

OpenStudy (owlcoffee):

I don't quite get the whole \[W(x)=lnx-lnW(x)\] How does that simplify the work?

OpenStudy (kainui):

https://repl.it/B3bI I'm busy helping out a friend right now, but that's what I'm talking about.

OpenStudy (kainui):

?? closed? :(

OpenStudy (whpalmer4):

Your program won't work for a value that gives a complex answer, though. Didn't know about repl.it, thanks!

OpenStudy (kainui):

It wasn't really supposed to work for complex answers, Java doesn't support complex values as a primitive type

OpenStudy (whpalmer4):

Yeah, was just pointing it out in case the OP decided to try using it to evaluate either the original problem or the one you suggested...